# Newton-Flow Particle Filters based on Generalized Cramér Distance

### v1

@@ IV-C Derivation of Recursive ODE
For deriving the recursive flow, we start with finite integration intervals [^3] This is different from directly taking discrete steps in $\gamma$ as in [11] as we integrate the ODE over the interval.. Imagine that we first integrate the ODE from $\gamma=0$ to some small $\gamma_{\ast}$ so that the posterior flows from the prior to the best permissible approximation $f(\underline{x},\underline{\eta}(\gamma_{\ast}))$ of the (impermissible) true posterior $\tilde{f}_{e}(\underline{x},\gamma_{\ast})\propto f_{p}(\underline{x})f_{L}(\underline{x},\gamma_{\ast})$. For DMD, the weights of the reference posterior have changed and are now not equal anymore. To keep it from degenerating further, we replace the original reference posterior $\tilde{f}_{e}(\gamma_{\ast})$ by its best approximation $f(\underline{x},\underline{\eta}(\gamma_{\ast}))$ at time $\gamma_{\ast}$. Form this time on, we compare with this newly reset reference density, that we then simply call reference posterior for the next steps. Of course, we have to modify the likelihood function that we use from $\gamma_{\ast}$ on as the new reference density already includes the likelihood integrated up to that time point. This involves calculating an effective likelihood $f_{L}^{\text{eff}}(\underline{x},\gamma)$ by dividing the original likelihood $f_{L}(\underline{x},\gamma)$ by the likelihood $f_{L}(\underline{x},\gamma_{\ast})$ used to far
$f_{L}^{\text{eff}}(\underline{x},\gamma)=\frac{f_{L}(\underline{x},\gamma)}{f_{L}(\underline{x},\gamma_{\ast})}\enspace.$ (26)
@@
$\dot{\underline{\eta}}(\gamma)=\underline{R}\big{(}\underline{\eta}(\gamma),f_{L}^{\text{eff}}(\underline{x},\gamma),\dot{f}_{L}^{\text{eff}}(\underline{x},\gamma)\big{)}\enspace.$ (27)
@@ V Generalized Cramér Distance
Instead of the general distance measure assumed in the previous section we now use the specific distance measure from the appendix that compares two DMDs that with (i) different numbers of components, (ii) at mutually distinct locations, (iii) with non-equal weights. For the particle flow in this paper, we use a specific configuration of the distance measure that compares two particle sets having the same numbers of samples.
@@ V-A Adaptation of General Distance Measure to Particle Flow
Let us define the given posterior density at artificial time $\gamma\in[0,1]$ as $\tilde{f}_{e}(\underline{x},\gamma)$. The tilde $\tilde{\phantom{\ast}}$ denotes the true posterior, but also the best approximation known at $\gamma$ when this is used as a stand-in for the true posterior. When $\tilde{f}_{e}(\underline{x},\gamma)$ represents a particle set, the dependency on $\gamma$ stems from its $\gamma$-dependent weights. This dependency is caused by the parametrized likelihood $f_{L}(\underline{x},\gamma)$ as defined in Sec. III used for weighting the samples in the Bayesian update step. From (8), by plugging in $f_{L}(\underline{x},\gamma)$, we obtain
@@
$\tilde{\underline{X}}=\begin{bmatrix}\tilde{\underline{x}}_{1},\tilde{\underline{x}}_{2},\ldots,\tilde{\underline{x}}_{L}\end{bmatrix}^{T}$ (29)
@@
$\tilde{\underline{w}}(\gamma)=\begin{bmatrix}\tilde{w}_{1}(\gamma),\tilde{w}_{2}(\gamma),\ldots,\tilde{w}_{L}(\gamma)\end{bmatrix}^{T}\enspace.$ (30)
@@
$\tilde{w}_{i}(\gamma)=\frac{1}{L}\frac{f_{L}(\tilde{\underline{x}}_{i},\gamma)}{F_{L}(\tilde{\underline{X}},\gamma)}$ (31)
@@
$F_{L}(\tilde{\underline{X}},\gamma)=\sum_{j=1}^{L}f_{L}(\tilde{\underline{x}}_{i},\gamma)\enspace.$ (32)
@@
$\underline{\eta}(\gamma)=\begin{bmatrix}\underline{x}_{1}(\gamma),\underline{x}_{2}(\gamma),\ldots,\underline{x}_{L}(\gamma)\end{bmatrix}^{T}$ (34)
@@
$\underline{w}=\begin{bmatrix}w_{1},w_{2},\ldots,w_{L}\end{bmatrix}^{T}\enspace.$ (35)
We define the distance $D$ between the true density $\tilde{f}_{e}(\underline{x},\gamma)$ and its approximation $f_{e}(\underline{x},\underline{w},\underline{\eta}(\gamma))$ based on the distance measure from the previous section as
$D(\underline{\eta}(\gamma),\gamma)=D\Bigl{(}\tilde{f}_{e}(\underline{x},\tilde{\underline{w}}(\gamma),\tilde{\underline{\eta}}),f_{e}\bigl{(}\underline{x},\underline{w},\underline{\eta}(\gamma)\bigr{)}\Bigr{)}\enspace.$ (36)
Again, we track the best approximation of the posterior that maintains equal weights starting with the posterior density for $\gamma=0$, i.e., $f_{e}\bigl{(}\underline{x},\underline{w},\underline{\eta}(0)\bigr{)}=f_{p}(\underline{x})$.
@@ V-B ODE for Iterative Particle Flow
During the particle flow, the densities $\tilde{f}_{e}$ and $f_{e}$ have the same number of components, so we have $M=L$. In that case, the gradient in (54) simplifies to
@@
$=$ $4\,w_{k}\Bigg{\{}\sum_{\begin{subarray}{c}i=1\\ i\neq k\end{subarray}}^{L}w_{i}\,(\underline{x}_{k}-\underline{x}_{i})\log\big{(}\|\underline{x}_{k}-\underline{x}_{i}\|^{2}\big{)}$
$\quad\quad-\sum_{i=1}^{L}\tilde{w}_{i}\,(\underline{x}_{k}-\tilde{\underline{x}}_{i})\log\big{(}\|\underline{x}_{k}-\tilde{\underline{x}}_{i}\|^{2}\big{)}\Bigg{\}}$
$+K_{2}\,w_{k}\sum_{i=1}^{L}(w_{i}\,\underline{x}_{i}-\tilde{w}_{i}\,\tilde{\underline{x}}_{i})$
With
$\operatorname{plog}(\underline{z})=\begin{cases}\underline{0}&\text{for }\underline{z}=\underline{0}\\ \underline{z}\,\log\left(\underline{z}^{T}\underline{z}\right)&\text{elsewhere}\end{cases}$ (38)
we can write
$\underline{G}_{k}=w_{k}\sum_{i=1}^{L}$ $\bigg{\{}4\left(w_{i}\,\operatorname{plog}\big{(}\underline{x}_{k}-\underline{x}_{i}\big{)}-\tilde{w}_{i}\,\operatorname{plog}\big{(}\underline{x}_{k}-\tilde{\underline{x}}_{i}\big{)}\right)$ (39)
$+K_{2}(w_{i}\,\underline{x}_{i}-\tilde{w}_{i}\,\tilde{\underline{x}}_{i})\bigg{\}}$
We now note that $\tilde{w}_{i}$ is a function of $\gamma$, i.e., $\tilde{w}_{i}=\tilde{w}_{i}(\gamma)$, while $\tilde{\underline{x}}_{i}$ and $w_{i}$ are independent of $\gamma$, and $\underline{\eta}_{i}$ only implicitly depends on $\gamma$. The derivative of the gradient vector w.r.t. $\gamma$ is denoted By
@@
$\underline{J}_{k}(\gamma)=$ $\frac{\partial G_{k}(\gamma)}{\partial\gamma}$ (41)
$=$ $-w_{k}\sum_{i=1}^{L}\bigg{\{}4\dot{\tilde{w}}_{i}(\gamma)\,\operatorname{plog}\big{(}\underline{x}_{k}-\tilde{\underline{x}}_{i}\big{)}+K_{2}\dot{\tilde{w}}_{i}(\gamma)\,\tilde{\underline{x}}_{i}\bigg{\}}\enspace.$
The normalized weights $\tilde{w}_{i}$ induced by the parametrized likelihood $f_{L}(\underline{x},\gamma)$ are given in (31) and (32). The derivatives of the normalized weights w.r.t. artificial time $\gamma$ are
$\dot{\tilde{w}}_{i}(\gamma)=\frac{1}{L}\frac{\dot{f}_{L}(\tilde{\underline{x}}_{i},\gamma)\,F_{L}({\mathbf{X}},\gamma)-f_{L}(\tilde{\underline{x}}_{i},\gamma)\,\dot{F}_{L}({\mathbf{X}},\gamma)}{F_{L}^{2}({\mathbf{X}},\gamma)}$ (42)
@@
$\dot{F}_{L}(\tilde{{\mathbf{X}}},\gamma)=\sum_{i=1}^{L}\dot{f}_{L}(\tilde{\underline{x}}_{i})\enspace.$ (43)
@@ V-C ODE for Recursive Particle Flow
$\underline{J}_{k}(\gamma)=-w_{k}\sum_{i=1}^{L}\bigg{\{}4\dot{w}_{i}(\gamma)\,\operatorname{plog}\big{(}\underline{x}_{k}-\underline{x}_{i}\big{)}+K_{2}\dot{w}_{i}(\gamma)\,\underline{x}_{i}\bigg{\}}\enspace.$ (44)
@@
$\dot{w}_{i}(\gamma)=\frac{1}{L}\bigg{\{}\frac{\dot{f}_{L}(\underline{x}_{i},\gamma)}{f_{L}(\underline{x}_{i},\gamma)}-\frac{1}{L}\sum_{j=1}^{L}\frac{\dot{f}_{L}(\underline{x}_{j},\gamma)}{f_{L}(\underline{x}_{j},\gamma)}\bigg{\}}\enspace.$ (45)
@@
The Hesse matrix is now also simplified. The diagonal block matrices are given by
${\mathbf{H}}_{kk}=K_{2}\,w_{k}^{2}\,{\mathbf{I}}_{N}$ (46)
@@ VI Conclusions
The proposed Newton-flow particle filter (NFPF) is a systematic approach to recursively estimate the hidden state of a nonlinear dynamic system. No heuristics are involved. The NFPF is easy to understand and simple to use as almost no tuning factors are required. It is fast for several reasons: (i) The number of particles to be processed is much smaller than in a standard particle filter. (ii) The distance measure, its gradient, and the Hesse matrix can be calculated in closed form. (iii) No optimization is required, just the solution of an ODE over a finite time interval. The proposed estimator shows a high performance compared to state-of-the-art approaches, produces high-quality estimates, is robust, is degeneration-free by design, and works in high dimensions. The proposed plugin replacement for a particle filter step is fully differentiable. This allows for end-to-end learning of the measurement nonlinearities or the noise densities, which will be exploited in future work.
@@
For convenience, we summarize the set distance measure derived in [7] and list all the required formulas in the slightly different notation used in this paper.
@@ VI-A Distance Measure
$\tilde{\underline{x}}_{j}=\begin{bmatrix}\tilde{\underline{x}}_{1,j},\tilde{\underline{x}}_{2,j},\ldots,\tilde{\underline{x}}_{N,j}\end{bmatrix}^{T}\in{\rm I\hskip-1.69998ptR}^{N}$ (47)
for $j=1,\ldots,M$. The approximating DMD has $L$ components with weights $w_{1},w_{2},\ldots,w_{L}$ and locations
@@
The Cramér-von Mises distance $D$ is given by
$D=D_{\tilde{x}\tilde{x}}-2D_{x\tilde{x}}+D_{xx}+K_{1}\,D_{E}\enspace,$ (49)
@@
$D_{\tilde{x}\tilde{x}}=\sum_{i=1}^{M}\sum_{j=1}^{M}\tilde{w}_{i}\,\tilde{w}_{j}\,\operatorname{xlog}\left(\|\tilde{\underline{x}}_{i}-\tilde{\underline{x}}_{j}\|_{2}^{2}\right)\enspace,$ (50)
$D_{x\tilde{x}}=\sum_{i=1}^{L}\sum_{j=1}^{M}w_{i}\,\tilde{w}_{j}\,\operatorname{xlog}\left(\|\underline{x}_{i}-\tilde{\underline{x}}_{j}\|_{2}^{2}\right)\enspace,$ (51)
$D_{xx}=\sum_{i=1}^{L}\sum_{j=1}^{L}w_{i}\,w_{j}\,\operatorname{xlog}\left(\|\underline{x}_{i}-\underline{x}_{j}\|_{2}^{2}\right)\enspace,$ (52)
and $\operatorname{xlog}(z)=z\cdot\log(z)$. $D_{E}$ can be viewed as a penalty term (with weight $c$) that ensures equal means and is given by
$D_{E}=\left\|\sum_{i=1}^{L}w_{i}\,\underline{x}_{i}-\sum_{j=1}^{M}\tilde{w}_{j}\,\tilde{\underline{x}}_{j}\right\|_{2}^{2}\enspace.$ (53)
$K_{1}$ is a large constant penalizing the squared difference between the means of $\tilde{f}_{e}(\underline{x})$ and $f_{e}(\underline{x})$.

### v2

@@ IV-C Derivation of Recursive ODE
For deriving the recursive flow, we start with finite integration intervals [^3] This is different from directly taking discrete steps in $\gamma$ as in [12] as we integrate the ODE over the interval.. Imagine that we first integrate the ODE from $\gamma=0$ to some small $\gamma_{\ast}$ so that the posterior flows from the prior to the best permissible approximation $f(\underline{x},\underline{\eta}(\gamma_{\ast}))$ of the (impermissible) true posterior $\tilde{f}_{e}(\underline{x},\gamma_{\ast})\propto f_{p}(\underline{x})f_{L}(\underline{x},\gamma_{\ast})$. For DMD, the weights of the reference posterior have changed and are now not equal anymore. To keep it from degenerating further, we replace the original reference posterior $\tilde{f}_{e}(\gamma_{\ast})$ by its best approximation $f(\underline{x},\underline{\eta}(\gamma_{\ast}))$ at time $\gamma_{\ast}$. From this time on, we compare with this newly reset reference density, that we then simply call reference posterior for the next steps. Of course, we have to modify the likelihood function that we use from $\gamma_{\ast}$ on as the new reference density already includes the likelihood integrated up to that time point. This involves calculating an effective likelihood $f_{L}^{\text{eff}}(\underline{x},\gamma)$ by dividing the original likelihood $f_{L}(\underline{x},\gamma)$ by the likelihood $f_{L}(\underline{x},\gamma_{\ast})$ used to far
$f_{L}^{\text{eff}}(\underline{x},\gamma)=\frac{f_{L}(\underline{x},\gamma)}{f_{L}(\underline{x},\gamma_{\ast})}\kern 5.0pt.$ (26)
@@
$\dot{\underline{\eta}}(\gamma)=\underline{R}\bigl(\underline{\eta}(\gamma),f_{L}^{\text{eff}}(\underline{x},\gamma),\dot{f}_{L}^{\text{eff}}(\underline{x},\gamma)\bigr)\kern 5.0pt.$ (27)
@@ V Generalized Cramér Distance
Instead of the abstract distance measure assumed in the previous section, we now use the specific distance measure in (48) from the Appendix that compares two DMDs that may have (i) different numbers of components $M$ and $L$, (ii) at mutually distinct locations, (iii) with non-equal weights. For the particle flow in this paper, we use a specific configuration of this distance measure that compares two particle sets having the same numbers of samples ($M=L$).
@@ V-A Adaptation of General Distance Measure to Particle Flow
Let us define the given posterior density at artificial time $\gamma\in[0,1]$ as $\tilde{f}_{e}(\underline{x},\gamma)$. When $\tilde{f}_{e}(\underline{x},\gamma)$ represents a particle set, the dependency on $\gamma$ stems from its $\gamma$-dependent weights. This dependency is caused by the parametrized likelihood $f_{L}(\underline{x},\gamma)$ as defined in Sec. III used for weighting the samples in the Bayesian update step. From (8), by plugging in $f_{L}(\underline{x},\gamma)$, we obtain
@@
$\tilde{\underline{X}}=\begin{bmatrix}\tilde{\underline{x}}_{1}^{T},\tilde{\underline{x}}_{2}^{T},\ldots,\tilde{\underline{x}}_{L}^{T}\end{bmatrix}^{T}$ (29)
@@
$\tilde{\underline{w}}(\gamma)=\begin{bmatrix}\tilde{w}_{1}(\gamma),\tilde{w}_{2}(\gamma),\ldots,\tilde{w}_{L}(\gamma)\end{bmatrix}^{T}\kern 5.0pt.$ (30)
@@
$\tilde{w}_{i}(\gamma)=\frac{f_{L}(\tilde{\underline{x}}_{i},\gamma)}{F_{L}(\tilde{\underline{X}},\gamma)}$ (31)
@@
$F_{L}(\tilde{\underline{X}},\gamma)=\sum_{i=1}^{L}f_{L}(\tilde{\underline{x}}_{i},\gamma)\kern 5.0pt.$ (32)
@@
$\underline{\eta}(\gamma)=\begin{bmatrix}\underline{x}_{1}^{T}(\gamma),\underline{x}_{2}^{T}(\gamma),\ldots,\underline{x}_{L}^{T}(\gamma)\end{bmatrix}^{T}$ (34)
@@
$\underline{w}=\begin{bmatrix}w_{1},w_{2},\ldots,w_{L}\end{bmatrix}^{T}\kern 5.0pt.$ (35)
We define the distance $D$ between the true density $\tilde{f}_{e}$ and its approximation $f_{e}$ based on the distance measure (48) as
$D(\underline{\eta}(\gamma),\gamma)=D\Bigl(\tilde{f}_{e}(\underline{x},\tilde{\underline{w}}(\gamma),\tilde{\underline{X}}),f_{e}\bigl(\underline{x},\underline{w},\underline{\eta}(\gamma)\bigr)\Bigr)\kern 5.0pt.$ (36)
The best approximation of the posterior that maintains equal weights is tracked as the minimum of $D$ over artificial time $\gamma$, starting with the posterior density for $\gamma=0$, i.e., $f_{e}\bigl(\underline{x},\underline{w},\underline{\eta}(0)\bigr)=f_{p}(\underline{x})$.
@@ V-B ODE for Iterative Particle Flow
During the particle flow, the densities $\tilde{f}_{e}$ and $f_{e}$ have the same number of components, so we have $M=L$. In that case, the gradient in (54) from the Appendix simplifies to
@@
$=$ $4\,w_{k}\Bigg\{\sum_{\begin{subarray}{c}i=1\\ i\neq k\end{subarray}}^{L}w_{i}\,\underline{\mathcal{G}}(\underline{x}_{k}-\underline{x}_{i})-\sum_{i=1}^{L}\tilde{w}_{i}\,\underline{\mathcal{G}}(\underline{x}_{k}-\tilde{\underline{x}}_{i})\Bigg\}$
$+2\,K\,w_{k}\sum_{i=1}^{L}(w_{i}\,\underline{x}_{i}-\tilde{w}_{i}\,\tilde{\underline{x}}_{i})$
We note that $\tilde{w}_{i}$ is a function of $\gamma$, i.e., $\tilde{w}_{i}=\tilde{w}_{i}(\gamma)$, while $\tilde{\underline{x}}_{i}$ and $w_{i}$ are independent of $\gamma$, and $\underline{\eta}_{i}$ only implicitly depends on $\gamma$. The derivative of the gradient vector w.r.t. $\gamma$ is given by
@@
$\underline{J}_{k}(\gamma)=$ $\frac{\partial\underline{G}_{k}(\gamma)}{\partial\gamma}$ (39)
$=$ $-w_{k}\sum_{i=1}^{L}\bigg\{4\,\dot{\tilde{w}}_{i}(\gamma)\,\underline{\mathcal{G}}(\underline{x}_{k}-\tilde{\underline{x}}_{i})+2\,K\dot{\tilde{w}}_{i}(\gamma)\,\tilde{\underline{x}}_{i}\bigg\}\kern 5.0pt.$
The normalized weights $\tilde{w}_{i}(\gamma)$ induced by the parametrized likelihood $f_{L}(\underline{x},\gamma)$ are given in (31) and (32). The derivatives of the normalized weights w.r.t. artificial time $\gamma$ are
$\dot{\tilde{w}}_{i}(\gamma)=\frac{\dot{f}_{L}(\tilde{\underline{x}}_{i},\gamma)\,F_{L}(\tilde{\underline{X}},\gamma)-f_{L}(\tilde{\underline{x}}_{i},\gamma)\,\dot{F}_{L}(\tilde{\underline{X}},\gamma)}{F_{L}^{2}(\tilde{\underline{X}},\gamma)}$ (40)
@@
$\dot{F}_{L}(\tilde{\underline{X}},\gamma)=\sum_{i=1}^{L}\dot{f}_{L}(\tilde{\underline{x}}_{i},\gamma)\kern 5.0pt.$ (41)
@@ V-C ODE for Recursive Particle Flow
$\underline{J}_{k}(\gamma)=-w_{k}\sum_{i=1}^{L}\bigg\{4\,\dot{w}_{i}(\gamma)\,\underline{\mathcal{G}}(\underline{x}_{k}-\underline{x}_{i})+2\,K\dot{w}_{i}(\gamma)\,\underline{x}_{i}\bigg\}\kern 5.0pt.$ (42)
@@
$\dot{w}_{i}(\gamma)=\frac{1}{L}\bigg\{\frac{\dot{f}_{L}(\underline{x}_{i},\gamma)}{f_{L}(\underline{x}_{i},\gamma)}-\frac{1}{L}\sum_{j=1}^{L}\frac{\dot{f}_{L}(\underline{x}_{j},\gamma)}{f_{L}(\underline{x}_{j},\gamma)}\bigg\}$ (43)
@@
The Hesse matrix in (56) from the Appendix is now also simplified. The diagonal block matrices are given by
${\mathbf{H}}_{kk}=2\,w_{k}^{2}\,\bigl\{K-2\,\log\bigl(d_{\text{min}}^{2}\bigr)\bigr\}\,{\mathbf{I}}_{N}$ (44)
@@ VI Conclusions
The proposed Newton-flow particle filter (NFPF) is a systematic approach to recursively estimate the hidden state of a nonlinear dynamic system. No heuristics are involved. The NFPF is easy to understand and simple to use as almost no tuning factors are required. It is fast for several reasons: (i) The number of particles to be processed is much smaller than in a standard particle filter. (ii) The distance measure, its gradient, and the Hesse matrix can be calculated in closed form. (iii) No optimization is required, just the solution of an ODE over a finite time interval. The proposed estimator shows a high performance compared to state-of-the-art approaches, produces unbiased high-quality estimates, is robust, is degeneration-free by design, and works in high dimensions. The proposed plugin replacement for a particle filter step is fully differentiable. This allows for end-to-end learning of the measurement nonlinearities or the noise densities, which will be exploited in future work.
@@
For convenience, we summarize the set distance measure derived in [7, 10] and list all the required formulas in the slightly different notation used in this paper.
@@ VI-A Distance Measure
$\tilde{\underline{x}}_{j}=\begin{bmatrix}\tilde{x}_{1,j},\tilde{x}_{2,j},\ldots,\tilde{x}_{N,j}\end{bmatrix}^{T}\in{\rm I\hskip-1.69998ptR}^{N}$ (46)
for $j=1,\ldots,M$. The approximating DMD $f(\underline{x})$ has $L$ components with weights $w_{1},w_{2},\ldots,w_{L}$ and locations
@@
The generalized Cramér-von Mises distance [7] is given by $\sqrt{D}$ with
$D=D\bigl(\tilde{f},f\bigr)=D_{\tilde{x}\tilde{x}}-2D_{x\tilde{x}}+D_{xx}+K\,D_{E}\kern 5.0pt,$ (48)
@@
$D_{\tilde{x}\tilde{x}}=\sum_{i=1}^{M}\sum_{j=1}^{M}\tilde{w}_{i}\,\tilde{w}_{j}\,\mathcal{D}(\tilde{\underline{x}}_{i}-\tilde{\underline{x}}_{j})\kern 5.0pt,$ (49)
$D_{x\tilde{x}}=\sum_{i=1}^{L}\sum_{j=1}^{M}w_{i}\,\tilde{w}_{j}\,\mathcal{D}(\underline{x}_{i}-\tilde{\underline{x}}_{j})\kern 5.0pt,$ (50)
$D_{xx}=\sum_{i=1}^{L}\sum_{j=1}^{L}w_{i}\,w_{j}\,\mathcal{D}(\underline{x}_{i}-\underline{x}_{j})\kern 5.0pt,$ (51)
and
$\mathcal{D}(\underline{d})=\underline{d}^{T}\underline{d}\,\log\bigl(d_{\text{min}}^{2}+\underline{d}^{T}\underline{d}\bigr)\kern 5.0pt.$ (52)
$D_{E}$ can be viewed as a penalty term that ensures equal means and is given by the squared difference between the means of $\tilde{f}(\underline{x})$ and $f(\underline{x})$
$D_{E}=\left\|\sum_{i=1}^{L}w_{i}\,\underline{x}_{i}-\sum_{j=1}^{M}\tilde{w}_{j}\,\tilde{\underline{x}}_{j}\right\|_{2}^{2}\kern 5.0pt.$ (53)
$K$ is a large constant for weighting the penalty term.
Please note that in contrast to [7], we added the small constant $d_{\text{min}}$ here to avoid singularities when two samples assume the same location.